. Using a hydrogen atom as an example, compare the gravitational 

and the electrostatic forces between a proton and an electron


1
Expert's answer
2021-05-03T18:44:20-0400

Using Coulomb’s Law


"F = \dfrac{k q_1 q_2}{ r^2}"


We have, "q_1 = q_2 = 1.6 \times 10^-{19} C"

"r = 0.05nm" (in the case of Hydrogen atom)


Hence,

"F = \dfrac{9 \times 10^9 \times 1.6 \times 10^{-19} \times 1.6 \times 10^{-19}}{(0.05\times 10^{-9})^2}"


"F = 9.216 \times 10^{-8}N"




Using Newton’s Universal Gravitational Law,


"F =\dfrac{ G m_1 m_2 }{ r^2}"


We can determine the strength of the gravitational force between the proton and electron.


Mass of electron "= m_1 = 9.1 \times 10^{-31}kg"


Mass of proton "= m_2 = 1.67 \times 10^{-27}kg"


Hence,

"F = \dfrac{6.67 \times10^{-11} \times 9.1 \times 10^{-31} \times 1.67 \times 10^{-27}}{(0.05 \times 10^{-9})^2}"


"F = 40.544 \times 10^{-45}N"


Hence, we can compare the gravitational and the electrostatic forces between a Proton and an Electron.